An axe that weighs 8.0 kg is used to strikes the log with a force of 400 N that comes to rest in 0.04
seconds. What is the magnitude of the impulse? How fast is the axe approach the log?
1
Expert's answer
2020-11-05T07:34:06-0500
"I=Ft=(400)(0.04)=16\ Ns"

"v=\frac{I}{m}=\frac{16}{8}=2.0\frac{m}{s}"